OpenStudy (trojanpoem):

After passing quantity of electricity in H2SO4 solution , 2 moles of H2 gas was formed, and when passing the same quantity of electricity in HCl Solution ..... moles of the same gas is formed. Explain your asnwer.

OpenStudy (aaronq):

Have you tried writing equations for this?

OpenStudy (aaronq):

Remember that electricity are just electrons, you're using these electrons to reduce \(H^+\) to \(H_2\).

OpenStudy (trojanpoem):

@aaronq , At first I was thinking of using Friday law : \[\frac{ Mas(x) }{ equivalentMass(x) } = \frac{ Mass(Y) }{ EM(Y) }\] But I don't have either equivalent masses.

OpenStudy (trojanpoem):

I think equivalent masses are both for (H) so Mass(X) = Mass(Y) Moles of X = Moles of Y = 2 ?

OpenStudy (aaronq):

you dont actually need any formulas, to answer the question to you only need to think conceptually and see where the electrons are going. You used the same amount of electricity (electrons) in both cases, right? ---- yes, 2 moles, the same amount of moles.

OpenStudy (trojanpoem):

Thanks. I never feel convinced of my answer until I assure it with formula. But was using the formula right ?

OpenStudy (trojanpoem):

And what was confusing my is that H2SO4 <- 2 H so in mole = 2 * NA HCl-> 1 H so in mole = 1 * NA

OpenStudy (aaronq):

yeah, but unless HCl was limited then you would get as much H2 as you could reduce. in other words: 2 HCl + e -> 2 H2 + Cl^-

OpenStudy (aaronq):

i guess using the formula works out in this case, although it's really not needed

OpenStudy (trojanpoem):

@aaronq , Thanks.

OpenStudy (aaronq):

no problems